# RD2Bench: Toward Data-Centric Automatic R&D

### v1

@@ Abstract
The progress of humanity is driven by those successful discoveries accompanied by countless failed experiments. Researchers often seek the potential research directions by reading and then verifying them through experiments. The process imposes a significant burden on researchers. In the past decade, the data-driven black-box deep learning method demonstrates its effectiveness in a wide range of real-world scenarios, which exacerbates the experimental burden of researchers and thus renders the potential successful discoveries veiled. Therefore, automating such a research and development (R&D) process is an urgent need. In this paper, we serve as the first effort to formalize the goal by proposing a Real-world Data-centric automatic R&D Benchmark, namely RD2Bench. RD2Bench benchmarks all the operations in data-centric automatic R&D (D-CARD) as a whole to navigate future work toward our goal directly. We focus on evaluating the interaction and synergistic effects of various model capabilities and aiding in selecting well-performing trustworthy models. Although RD2Bench is very challenging to the state-of-the-art (SOTA) large language model (LLM) named GPT-4, indicating ample research opportunities and more research efforts, LLMs possess promising potential to bring more significant development to D-CARD: They are able to implement some simple methods without adopting any additional techniques. We appeal to future work to take developing techniques for tackling automatic R&D into consideration, thus bringing the opportunities of the potential revolutionary upgrade to human productivity.
@@ 1 Introduction
The advancement of human society and the enhancement of living standards are highly correlated with the development of technology (Smith, 1937; Ranis & Fei, 1961; Perez, 2003; Brynjolfsson & McAfee, 2014). Numerous truths and principles remain undiscovered in the world, awaiting experimental exploration (Shapere, 1964; Popper, 2005). Those few successful discoveries, accompanied by countless failed experiments, propel the frontiers of technology. Historically, scientific researchers, including Edison, have undertaken extensive experiments by conducting them manually. In the age of AI, the influence of data-driven solutions, such as machine learning (ML) systems, is rapidly expanding (Mikolov et al., 2013; Devlin et al., 2018; OpenAI, 2023b). These systems are known for their robust fitting capabilities and their “black box” nature, which significantly increases the experimental load on researchers and hinders the process of identifying and validating effective methodologies. This paper concentrates on this critical scenario, which we refer to as Data-Centric Research and Development (R&D). To cope with the prohibitively expensive costs and the overwhelming volume of experiments required, we consider automating such an R&D process for higher research efficiency by leveraging the strong language understanding and programming ability of the state-of-the-art (SOTA) large language models (LLMs) (Srivastava et al., 2023). The brief illustration of Data-Centric Automatic R&D (D-CARD) is shown in Figure 1.
@@
Figure 1: An overview of the R&D process. Researchers read papers and reports to extract the implementable methods (usually formulated as mathematical formulas or model architectures) for seeking potential research directions. Then, they correctly implement the methods to obtain the results for further analysis and development.
@@
The first step towards automatic R&D is to formalize the task and provide a benchmark for identifying the potential effective methods and research directions. Intuitively, an outstanding methodology identified by the benchmark should possess (1) strong language understanding ability to identify the implementable methods or ideas (e.g., formulations and models) in the given raw information (e.g., papers, reports, websites, etc.) and (2) strong implementation ability to accurately implement the methods by programming and then obtain reliable experimental results. Previous work focuses on benchmarking the different aspects of the two abilities. Specifically, the language understanding ability of LLMs is partly evaluated through analyzing their performance on relation extraction (Wadhwa et al., 2023), question answering (Zhuang et al., 2023), and other natural language processing (NLP) tasks (Qin et al., 2023a). Meanwhile, the implementation ability of LLMs is partly tested through the benchmarks like SWE-Bench (Jimenez et al., 2023b), ToolBench (Qin et al., 2023c), ML-Bench (Liu et al., 2023b) and MetaTool (Anonymous, 2024), which study their ability of solving GitHub issues, using tools to program, and determining whether to use tools in a given scenario.
@@
In this paper, we serve as the first effort to investigate the capabilities of the SOTA LLMs in tackling automatic R&D and propose a Real-world Data-centric automatic R&D Benchmark (RD2Bench). The scenario studied by RD2Bench possesses two unique and distinct characteristics that fundamentally differentiate it from others. First, RD2Bench focuses on studying the real-world scenario where all the operations in R&D are automatic and evaluated as a whole, thus navigating the related future research efforts toward the goal of developing human technology more effectively. The real-world scenario requires more comprehensive and advanced model capabilities and exhibits new challenges. Second, we study the real-world automatic R&D in data-centric settings to navigate future work toward the urgent experimental exploration need brought by black-box data-driven models. Compared with existing benchmarks, RD2Bench possesses two significant advantages:
@@
(1) RD2Bench evaluates the interaction and synergistic effects of various model capabilities instead of focusing on a single aspect of ability, which not only captures the frontier of SOTA LLMs but also bridges the gap between studying “individual ability” and “real-world synergistic effects of abilities”. In automatic R&D, an ML system fails to complete the task even if it possesses both the strong information extraction ability and the strong programming or tool-using ability: While it succeeds in extracting methods and implementing them, it fails in selecting the appropriate data from the datasets or misunderstanding either the descriptions of data features or the requirements expressed by prompts. Additionally, exhaustively enumerating all the aspects for benchmarking is extremely challenging, which is overcame by RD2Bench.
@@
(2) RD2Bench tends to select well-performing trustworthy models instead of those models that fail to learn rationales and causality yet possess outstanding performance. Specifically, ML systems easily achieve SOTA performance on previous benchmarks by shortcut learning or learning spurious correlations instead of learning rationales or causality (Mudrakarta et al., 2018; Geirhos et al., 2020; Cui & Athey, 2022; Wang et al., 2022; Chen et al., 2023). This renders a benchmark ineffective and misleading as it fails to accurately identify the well-performing trustworthy methods. For example, an ML system achieves SOTA performance on dog classification by merely recognizing grass (Zhang et al., 2021). RD2Bench, on the contrary, eliminates such models by its high difficulty and large scope. The decision rules of models have to simultaneously satisfy at least four major requirements: (1) accurately and comprehensively extracting the implementable methods; (2) precisely selecting the method-specific data for computation; (3) correctly writing the code according to the logic expressed by methods and prompts; (4) successfully storing the correct results in a predefined format. Therefore, the decision rules of models selected by this benchmark are stable (work well in various situations), and thus getting closer to rationales and causality (Cui & Athey, 2022).
@@
We evaluate the existing SOTA LLMs on RD2Bench to expose the bottleneck of them and characterize the future research direction. RD2Bench reveals new insights: (1) Among the popular LLMs, GPT-4 exhibits promising potency in dealing with the D-CARD task; (2) Detailed information of data descriptions significantly improves the performance of GPT-4; (3) The ability to query domain-specific knowledge is a basic requirement of D-CARD methods; (4) The more complex the method is, the more unstable the model performance is.
2 RD2Bench
@@
Overall, our benchmark focuses on evaluating the finally implemented results according to the given raw information (e.g., papers, reports, websites, etc.). Moreover, we also provide human-annotated ground-truth information corresponding to the intermediate steps for debugging and more comprehensive evaluation, which not only aids in selecting the most trustworthy models by tracing back model operations in each step but also facilitates the research efficiency by clearly locating the problems. RD2Bench selects those well-performing trustworthy models that are able to accurately calculate the final results with precise operation in each step. We introduce the details of our proposed RD2Bench in the following sections. In section 2.1, we introduce the details of human-annotated ground truth and how we collect data to form RD2Bench. Then, we elaborate on the two necessary steps, namely method extraction and method implementation, to perform R&D in section 2.2 and section 2.3. Finally, we detail our adopted metrics in section 2.4.
@@ 2.1 Data Collection
Figure 2: An example of formula implementation task.
As mentioned in the previous section, we are aiming to study the data-centric automatic R&D tasks and take formula implementation and model architecture implementation as the two major tasks included in this benchmark.
For formula implementation, we collect finance reports and stock trading data as real-world input. More specifically, the formula implementation task is to implement the factors function (usually mathematical formulas that take complex numeric input data about stock, company, and market as input and output a series of values with the time series) proposed in the report. To be more comprehensive, we include 11 factors from three major categories (fundamental, price-volume, high-frequency, denoted as Data I, II, III in following) and four difficulty levels (easy, medium, hard, and new factor). The difficulty level is manually labeled by domain experts according to the complexity of implementation (code implementation and idea understanding) and the novelty of the task. All ground truth implementations and formulas of the factors function are provided by us in the dataset. With the help of domain experts, the ground truth implementations and formulas are reliable and clear. One example of the formula implementation task is shown in Figure 2.
For model architecture implementation, we collect papers including  (Gravina et al., 2023; Rossi et al., 2023; Rampášek et al., 2022; Lim et al., 2021; Yang et al., 2023a; Wang et al., 2024) and corresponding ground truth implementation codes using pytorch (Paszke et al., 2019) and torch_gemometric framework (Fey & Lenssen, 2019) in deep learning, more specifically, in graph representation learning field. The model architecture implementation task is to implement the new model architecture or layer proposed in the paper. We also manually label the difficulty level (easy, medium, hard) of the task based on the complexity of implementation (computational graphs and tensor operations) and the novelty of the task (brand new mechanism or new structure). All the ground truth implementations are provided by excellent researchers and engineers, which are reliable and truthful for evaluation. See the  A.2 for more details about the dataset and the task.
2.2 Method Extraction
@@
This section evaluates the model’s ability to recognize and extract information from formulas and to innovate within an automated R&D context. The model must identify actionable methods from extensive research data and extract all necessary information to realize these methods, which is critical for the subsequent code implementation.
This initial benchmark session tests the model’s capability to accurately and comprehensively extract all conditions required for the methods it analyzes, and to develop the corresponding code implementation. The model should not proceed with incomplete methods but must fully understand and code for complete ones. We specify a format for information extraction that aligns with a manually labeled format or an appendix. We measure the model’s extraction accuracy using the F1 score, considering both entity recognition and mention. For description accuracy, we compare the model’s output with a manually annotated template, aiming for over 90% similarity. Developers can manually review this for precise evaluation. The model may encounter methods in the source material that are mentioned by name without detailed formulas or definitions. In such cases, the model can opt not to extract information or infer details based on the material’s semantics. We encourage the latter approach to foster creativity and support future research that may introduce new methods and formulas. Currently, only methods explicitly named in the material are subject to this creative review, but future benchmarks may assess the model’s ability to introduce entirely new concepts without prior mentions.
As most of our input data is encoded in the form of document files, we first use parsing tools to extract text content from files. Azure document intelligence API (4.0) is used for parsing reports and academic papers in PDF format.
2.3 Method Implementation
@@
In this section, we evaluate the performance of LLM in the implementation of methods. Given all the necessary conditions provided to the model after the previous step, the model needs to select the necessary data and write code from scratch to implement the method with an informative and well-organized prompt. Details of the prompt are included in the dataset, which is also shown in  A.3. We encourage model use of Python and popular data analysis and machine learning libraries.
2.4 Metrics
@@
The goal of the benchmark is to evaluate the performance of LLM in data-centric R&D tasks, correctness and efficiency are the two major aspects we care about. To be more comprehensive and reveal more signals of performance, we propose multiple metrics for each task and each stage of the task. For formula implementation, we care about the average and maxima “running success rate”, “format success rate”, “Pearson correlation” and “value accuracy” across multiple independent attempts. To be simple, “avg.”, “exe.”, “form.”, “corr.”, “acc.” denotes the average value, number of successful execution times, number of the matched result formats, the correlation, and the accuracy of corresponding values, respectively. See more details about the metrics calculation details in App  A.1.
@@
For model architecture implementation, we believe a successful implementation of a model should be consistent with the ground truth implementation as the model be viewed as a numeric function and combination of tensor transformations. Therefore, we propose these two metrics for the model architecture implementation task: tensor shape consistency rate (tsc.), tensor value consistency rate (tvc.). To be more specific, for each layer of the model, we calculate the consistency rate of the tensor shape and tensor value between the ground truth implementation and the implementation generated by the LLM. All the ground truth tensor value is determined by ground truth implementation codes with random Gaussian noise. Therefore, the formula for the two metrics is as follows, where $S_{\text{shape}}^{i}$ and $S_{\text{value}}^{i}$ are the consistency rate of tensor shape and tensor value in layer $i$, respectively, and $d_{i}$ is the maximum length of the two tensors as the two tensors are $\mathbf{Z}_{i}$ and $\mathbf{Z}_{i}^{*}$, the ground truth and the generated tensor, respectively:
@@
$S_{\text{shape}}^{i}(\mathbf{Z}_{i},\mathbf{Z}_{i}^{*})$ $=\left(1+\exp\left(\frac{\sum_{j=1}^{d}|\text{dim}(\mathbf{Z}_{i})_{j}-\text{dim}(\mathbf{Z}_{i}^{*})_{j}|}{d}\right)\right)^{-1},$ (1)
$S_{\text{value}}^{i}(\mathbf{Z}_{i},\mathbf{Z}_{i}^{*})$ $=\left(1+\exp\left(\frac{\sum_{j=1}^{d}|\mathbf{Z}_{i}^{(j)}-\mathbf{Z}_{i}^{*(j)}|}{d}\right)\right)^{-1},$
$d$ $=\max(\text{len}(\text{dim}(\mathbf{Z}_{i})),\text{len}(\text{dim}(\mathbf{Z}_{i}^{*}))),$
@@
while the shorter tensor is padded with zeros to match the length of the longer tensor. As the final score of the two metrics, we use the weighted sum of the consistency rate of all layers, weight increases with the depth of the layer and is summed as one.
$S_{\text{final}}=\frac{\sum_{i=1}^{n}S^{i}\cdot\gamma^{i}}{\sum_{i=1}^{n}\gamma^{i}},$ (2)
where $n$ is the number of layers in the model, $\gamma$ is a tunable hyperparameter to control the weight increase, and we set $\gamma=1.1$ in our experiments.
An example of the calculation is shown in Figure 3, using model LinkX  (Lim et al., 2021) as an example. Meanwhile, we also include ”average running success rate” as the basic metric for the model architecture implementation task, which is the same as the formula implementation task.

### v2

# Towards Data-Centric Automatic R&D

@@ Abstract
The progress of humanity is driven by those successful discoveries accompanied by countless failed experiments. Researchers often seek the potential research directions by reading and then verifying them through experiments. The process imposes a significant burden on researchers. In the past decade, the data-driven black-box deep learning method has demonstrated its effectiveness in a wide range of real-world scenarios, which exacerbates the experimental burden of researchers and thus renders the potential successful discoveries veiled. Therefore, automating such a research and development (R&D) process is an urgent need. In this paper, we serve as the first effort to formalize the goal by proposing a Real-world Data-centric automatic R&D Benchmark, namely $\text{RD}^{2}$ench. $\text{RD}^{2}$ench benchmarks all the operations in data-centric automatic R&D (D-CARD) as a whole to navigate future work toward our goal directly. We focus on evaluating the interaction and synergistic effects of various model capabilities and aiding in selecting well-performing trustworthy models. Although $\text{RD}^{2}$ench is very challenging to the state-of-the-art (SOTA) large language model (LLM) named GPT-4, indicating ample research opportunities and more research efforts, LLMs possess promising potential to bring more significant development to D-CARD: They are able to implement some simple methods without adopting any additional techniques. We appeal to future work to take developing techniques for tackling automatic R&D into consideration, thus bringing the opportunities of the potential revolutionary upgrade to human productivity.
@@ 1 Introduction
footnotetext: This is an open-source project starting in Oct. 2023. 1 Work done during an internship at Microsoft.
@@
The advancement of human society and the enhancement of living standards are highly correlated with the development of technology [39, 33, 26, 4]. Numerous truths and principles remain undiscovered in the world, awaiting experimental exploration [36, 27]. Those few successful discoveries, accompanied by countless failed experiments, propel the frontiers of technology. Historically, scientific researchers, including Edison, have undertaken extensive experiments by conducting them manually. In the age of AI, the influence of data-driven solutions, such as machine learning (ML) systems, is rapidly expanding [19, 7, 21]. These systems are known for their robust fitting capabilities and their “black box” nature, which significantly increases the experimental load on researchers and hinders the process of identifying and validating effective methodologies. This paper concentrates on this critical scenario, which we refer to as Data-Centric Research and Development (R&D). To cope with the prohibitively expensive costs and the overwhelming volume of experiments required, we consider automating such an R&D process for higher research efficiency by leveraging the strong language understanding and programming ability of the state-of-the-art (SOTA) large language models (LLMs) [40]. The brief illustration of Data-Centric Automatic R&D (D-CARD) is shown in Figure 1.
@@
Figure 1: An overview of the R&D process. Researchers read papers and reports to extract the implementable methods (usually formulated as mathematical formulas or model architectures) for seeking potential research directions. Then, they accurately implement the methods to obtain the results for further analysis and development.
@@
The first step towards automatic R&D is to formalize the task and provide a benchmark for identifying the potential effective methods and research directions. Intuitively, an outstanding methodology identified by the benchmark should possess (1) strong language understanding ability to identify the implementable methods or ideas (e.g., formulations and models) in the given raw information (e.g., papers, reports, websites, etc.) and (2) strong implementation ability to accurately implement the methods by programming and then obtain reliable experimental results. Previous work focuses on benchmarking the different aspects of the two abilities. Specifically, the language understanding ability of LLMs is partly evaluated through analyzing their performance on relation extraction [43], question answering [57], and other natural language processing (NLP) tasks [29]. Meanwhile, the implementation ability of LLMs is partly tested through benchmarks like SWE-Bench [11], ToolBench [30], ML-Bench [16] and MetaTool [1], which study their ability of solving GitHub issues, using tools to program, and determining whether to use tools in a given scenario.
@@
In this paper, we serve as the first effort to investigate the capabilities of the SOTA LLMs in tackling automatic R&D and propose a Real-world Data-centric automatic R&D Benchmark ($\text{RD}^{2}$ench). The scenario studied by $\text{RD}^{2}$ench possesses two unique and distinct characteristics that fundamentally differentiate it from others. First, $\text{RD}^{2}$ench focuses on studying the real-world scenario where all the operations in R&D are automatic and evaluated as a whole, thus navigating the related future research efforts toward the goal of developing human technology more effectively. The real-world scenario requires more comprehensive and advanced model capabilities and exhibits new challenges. Second, we study the real-world automatic R&D in data-centric settings to navigate future work toward the urgent experimental exploration need brought by black-box data-driven models. Compared with existing benchmarks, $\text{RD}^{2}$ench possesses two significant advantages:
@@
(1) $\text{RD}^{2}$ench evaluates the interaction and synergistic effects of various model capabilities instead of focusing on a single aspect of ability, which not only captures the frontier of SOTA LLMs but also bridges the gap between studying “individual ability” and “real-world synergistic effects of abilities”. In automatic R&D, an ML system fails to complete the task even if it possesses both the strong information extraction ability and the strong programming or tool-using ability: While it succeeds in extracting methods and implementing them, it fails in selecting the appropriate data from the datasets or misunderstanding either the descriptions of data features or the requirements expressed by prompts. Additionally, exhaustively enumerating all the aspects for benchmarking is extremely challenging, which is overcome by $\text{RD}^{2}$ench.
@@
(2) $\text{RD}^{2}$ench tends to select well-performing trustworthy models instead of those models that fail to learn rationales and causality yet possess outstanding performance. Specifically, ML systems easily achieve SOTA performance on previous benchmarks by shortcut learning or learning spurious correlations instead of learning rationales or causality [20, 9, 6, 46, 5]. This renders a benchmark ineffective and misleading as it fails to accurately identify the well-performing trustworthy methods. For example, an ML system achieves SOTA performance on dog classification by merely recognizing grass [55]. $\text{RD}^{2}$ench, on the contrary, eliminates such models by its high difficulty and large scope. The decision rules of models have to simultaneously satisfy at least four major requirements: (1) accurately and comprehensively extracting the implementable methods; (2) precisely selecting the method-specific data for computation; (3) correctly writing the code according to the logic expressed by methods and prompts; (4) successfully storing the correct results in a predefined format. Therefore, the decision rules of models selected by this benchmark are stable (work well in various situations), and thus getting closer to rationales and causality [6].
@@
We evaluate the existing SOTA LLMs on $\text{RD}^{2}$ench to expose their bottleneck and characterize the future research direction. $\text{RD}^{2}$ench reveals new insights: (1) Among the popular LLMs, GPT-4 exhibits promising potency in dealing with the D-CARD task; (2) Detailed information of data descriptions significantly improves the performance of GPT-4; (3) The ability to query domain-specific knowledge is a basic requirement of D-CARD methods; (4) The more complex the method is, the more unstable the model’s performance is.
2 Related Work
2.1 LLM as Autonomous Agent
In the past few years, LLM has made great achievements in both academia and industry [22, 42], and has achieved results that surpass the previous level in a number of classic tasks [56]. Research has shown that with the growth of data volume and model size [58], LLM has emerged with stronger reasoning and other capabilities [23]. These capabilities enable LLM to exhibit certain agent-like behaviors in some tasks such as using or creating tools [31, 28], planning [54, 3], and memory. Therefore, more and more researchers have expressed their expectations for its human-like and overall capabilities, and have made preliminary explorations of it as an independent agent [44, 38]. Multi-agent collaboration [49, 14] is also introduced to LLM for better accuracy and generalizability. Moreover, for reducing human efforts and automatically exploring, previous work focuses on autonomous LLM agents for general purpose are purposed [51, 37]. Positive views further believe that the realization of AGI may come from the evolution of autonomous LLM and some inspiring examples have been released [25].
However, most research still focuses on limited scenarios that are given with clear and fixed questions and backgrounds. A recent work [53] has attempted to introduce LLM to the R&D field and formalize the R&D process as a sequence of tasks. However, there is no easy-to-use benchmark for the community and current R&D tasks may be too general and can’t reveal significant signals. In this work, we propose a benchmark for LLM in data-centric R&D tasks and provide a comprehensive evaluation.
2.2 Semi-Automatic R&D with Agents
Scientific research and development (R&D) is a time-consuming and important process. In the past, R&D has been mainly conducted by human researchers with countless failed experimental explorations and creative observation conclusions. Agents have been introduced to R&D to reduce human efforts and automatically explore. Recently, there have been attempts to partly automate R&D, including the automatic chemical synthesis planning  [2], automatic molecular design  [13, 35, 2], automatic theorem proving  [45, 52]. However, these attempts mainly focus on automatic searching for possible solutions and optimizations with symbolic representation  [18] and heuristic techniques  [48], but less addressing long-horizon planning, implementation, and reasoning for the next step idea exploration. Moreover, the data-centric R&D tasks currently have not been explored in the community, and no benchmark has been proposed for the community. Previous works have applied LLM to real-world R&D tasks such as debugging issues  [41, 12] or only focus on data-centric but not real-world R&D tasks  [17]. In this work, we propose a benchmark for LLMs in data-centric R&D tasks and evaluate the performance of LLMs.
3 $\text{RD}^{2}$ench
@@
Overall, our benchmark focuses on evaluating the finally implemented results according to the given raw information (e.g., papers, reports, websites, etc.). Moreover, we also provide human-annotated ground-truth information corresponding to the intermediate steps for debugging and more comprehensive evaluation. $\text{RD}^{2}$ench selects well-performing models that follow human operations and accurately calculate the final results. We introduce the details of our proposed $\text{RD}^{2}$ench in the following sections. In section 3.1 and section 3.2, we introduce how we collect data and perform human annotation to form $\text{RD}^{2}$ench. Then, we elaborate on the two necessary steps, namely method extraction and method implementation, to perform R&D in section 3.3 and section 3.4. Finally, we detail our adopted evaluation metrics in section 3.5.
@@ 3.1 Data Collection
We consider the raw information that contains formulas and models, which represent a wide range of methods proposed in the AI domain.
Data Collection with Formulas. We prepare raw information that contains formulas as the input of R&D. Raw information is presented as publicly available financial reports and stock trading data. Formulas are usually mathematical formulas that take complex numeric input data about stock, company, and market as input and output a series of values with the time series. We collect financial reports with 27 implementable formulas distributed in three difficulty levels: easy, medium, and hard. Domain experts manually label the difficulty levels according to the complexity of implementation. To obtain their implementation results, an agent is expected to accurately select the features from three types of trading data scattered across 2010 to 2022, namely fundamental, price-volume, and high-frequency data. We denote the three types of data as Data I, II, and III, respectively.
Data Collection with Models. We collect papers with six open-sourced models [10, 34, 32, 15, 50, 47]. The implementation of models adopts Pytorch [24] and torch_gemometric framework [8] to perform deep learning. All the papers and models are publicly available. We manually label the difficulty level (easy, medium, hard) of the task based on the complexity of implementation (computational graphs and tensor operations). We refer the readers to the appendix for more details about the dataset and the task.
3.2 Human Annotation.
To provide a more comprehensive evaluation for debugging and analyzing, we conduct human annotation to provide the ground-truth results of our collected data, namely method extraction results and method implementation results.
Challenges. We confront five main challenges in the human annotation process. First, we need to identify the difficulty levels of methods to ensure the diversity of our benchmark and expose the bottleneck of current models. Second, we have to identify and discard the raw information if its presented methods demand unavailable data: The computation of some formulas can require confidential information that is not publicly available. Third, since the definitions or descriptions of some methods can be vague, leading them to be unimplementable, we have to filter out these methods. Fourth, some domain-specific methods containing factual errors should be filtered out since they are not implementable. Fifth, we should distinguish the domains and types of the methods according to their descriptions. To sum up, all the challenges imply the fact that human annotation of $\text{RD}^{2}$ench requires high time cost and expertise of annotators. Therefore, we commit more effort to designing the annotation guidelines and mechanisms to ensure the dataset quality.
Annotation Guidelines and Mechanisms. We provide principled guidelines and mechanisms to the annotators. Specifically, we first clarify our requirements in the guidelines: The extracted ground-truth methods are implementable without the requirement of any additional information, and the ground-truth implementation results are stored in an expected space with the predefined form and shape. Second, we provide training to the annotators. After the training step, we delicately design a test case and an interview to examine if the annotators possess expertise in the target domain (e.g., financial domain) and understand the guidelines. Third, we make each annotation result checked at least twice by annotators. A senior researcher and an applied scientist with domain expertise and rich engineering experience will further check each annotation result together before it finally gets included in the benchmark.
3.3 Method Extraction Step
@@
We evaluate the ability of models to recognize and extract information from raw information (e.g., R&D context). A qualified model is expected to discern feasible methods (formulas and models) from extensive research materials and extract all necessary information for implementing these formulas. The ability serves as the foundational premise for subsequent code implementation.
We expect the model to accurately and comprehensively extract the methods mentioned in the research materials it reviews, including all essential conditions required to implement the method. For methods with incomplete information, further implementation is not required; for methods with complete conditions, a model is expected to correctly comprehend the semantic meaning of these conditions stated in natural language and generate corresponding code. Specifically, we have predefined the extraction format (key-value pair) for the model. We employ the F1 score to measure the comprehensiveness and accuracy of method identification and extraction.
Note that some methods in the original materials might only imply their function, effect, or origin through their names without explicitly presenting their formulas, definitions, or other details. In such cases, the model may choose not to extract them or opt to autonomously complete them based on the semantics of the original materials. We expect the latter approach in future work, as it showcases the creativity of models by proposing new formulas and generating brand-new, informative, and reliable information. In the current version of the benchmark, only methods mentioned by name are evaluated in this manner; future iterations will explore and assess the model’s ability to generate new names and formulas when none are explicitly mentioned.
3.4 Method Implementation Step
@@
In this section, we evaluate the performance of LLM in the implementation of methods. Given all the necessary conditions provided to the model after the previous step, the model needs to select the necessary data and write code from scratch to implement the method with an informative and well-organized prompt. Details of the prompt are included in the dataset, which is also shown in the appendix. We encourage models to use Python and perform data analysis. They are also permitted to use common machine-learning libraries. One example of the method implementation step is shown in Figure 2.
Figure 2: An example of a formula implementation task.
3.5 Evaluation Metrics
@@
We adopt multiple metrics to evaluate model performance in each step. For formula implementation, we adopt the average and maxima “running success rate”, “format success rate”, “Pearson correlation” and “value accuracy” across multiple independent attempts. We use “avg.”, “exe.”, “form.”, “corr.”, and “acc.” to denote the average value, number of successful execution times, number of matched result formats, the correlation, and the accuracy of corresponding values, respectively. We refer the readers to more details about the metrics calculation details in App A.
@@
For model implementation, we believe a successful implementation of a model should be consistent with the ground truth implementation as the model can be viewed as a numeric function and combination of tensor transformations. Therefore, we propose these two metrics for the model architecture implementation task: tensor shape consistency rate (tsc.), tensor value consistency rate (tvc.). Specifically, for each model layer, we calculate the consistency rate of the tensor shape and tensor value between the ground truth implementation and the implementation generated by the LLM. All the ground truth tensor values are determined by ground truth implementation codes with random Gaussian noise. Therefore, the formula for the two metrics is as follows, where $S_{\text{shape}}^{i}$ and $S_{\text{value}}^{i}$ are the consistency rate of tensor shape and tensor value in layer $i$, respectively, and $d_{i}$ is the maximum length of the two tensors as the two tensors are $\mathbf{Z}_{i}$ and $\mathbf{Z}_{i}^{*}$, the ground truth and the generated tensor, respectively:
@@
$\begin{gathered}S_{\text{shape}}^{i}(\mathbf{Z}_{i},\mathbf{Z}_{i}^{*})=\left(1+\exp\left(\frac{\sum_{j=1}^{d}|\text{dim}(\mathbf{Z}_{i})_{j}-\text{dim}(\mathbf{Z}_{i}^{*})_{j}|}{d}\right)\right)^{-1},\\ S_{\text{value}}^{i}(\mathbf{Z}_{i},\mathbf{Z}_{i}^{*})=\left(1+\exp\left(\frac{\sum_{j=1}^{d}|\mathbf{Z}_{i}^{(j)}-\mathbf{Z}_{i}^{*(j)}|}{d}\right)\right)^{-1},d=\max(\text{len}(\text{dim}(\mathbf{Z}_{i})),\text{len}(\text{dim}(\mathbf{Z}_{i}^{*}))),\end{gathered}$ (1)
@@
while the shorter tensor is padded with zeros to match the length of the longer tensor. As the final score of the two metrics, we use the weighted sum of the consistency rate of all layers, weight increases with the depth of the layer and is summed as one: $S_{\text{final}}=\frac{\sum_{i=1}^{n}S^{i}\cdot\gamma^{i}}{\sum_{i=1}^{n}\gamma^{i}},$ where $n$ is the number of layers in the model, $\gamma$ is a tunable hyperparameter to control the weight increase, and we set $\gamma=1.1$ in our experiments.
@@
An example of the calculation is shown in Figure 3, using model LinkX  [15] as an example. Meanwhile, we also include the “average running success rate” as the basic metric for the model architecture implementation task, which is the same as the formula implementation task.